# Inferring Scientific Cross-Document Coreference and Hierarchy with Definition-Augmented Relational Reasoning

### v2

@@ Abstract
We address the fundamental task of inferring cross-document coreference and hierarchy in scientific texts, which has important applications in knowledge graph construction, search, recommendation and discovery. LLMs can struggle when faced with many long-tail technical concepts with nuanced variations. We present a novel method which generates context-dependent definitions of concept mentions by retrieving full-text literature, and uses the definitions to enhance detection of cross-document relations. We further generate relational definitions, which describe how two concept mentions are related or different, and design an efficient re-ranking approach to address the combinatorial explosion involved in inferring links across papers. In both fine-tuning and in-context learning settings we achieve large gains in performance. We provide analysis of generated definitions, shedding light on the relational reasoning ability of LLMs over fine-grained scientific concepts.111We make code, models and generated definitions available https://github.com/TomHopeLab/ScicoRadar.
@@ 1 Introduction
As the volume of scientific papers continues to explode, so does the intricate web of concepts mentioned across papers. Papers in computer science, for example, discuss fine-grained concepts such as algorithms, methods, tasks and problem areas Luan et al. (2018). Being able to automatically identify, e.g., when papers are referring to the same problem while using very different language (coreference), or that a method used in one paper is part of a more general family of methods (hierarchy)—could help unlock a wide range of applications, including in scientific knowledge-base construction Mondal et al. (2021); Hope et al. (2021), information retrieval Wang et al. (2023); Kang et al. (2024); Wu et al. (2024), recommendation systems Mysore et al. (2023), systems for paper reading and learning new topics Murthy et al. (2022); Guo et al. (2024), and even in computational creativity for science Hope et al. (2017); Portenoy et al. (2022); Wang et al. (2024); Srinivasan and Chan (2024).
@@
Detecting cross-document coreference and hierarchy relations in scientific papers poses significant challenges Cattan et al. (2021b). Concepts manifest in long-tail, nuanced variations. Technical language is characterized by ambiguity, diversity and fine-grained levels of concept abstractions. For example, energy-based network embedding falls under neural representation learning despite having different surface forms; while low energy embedded networks in communications is not related to energy-based network embedding despite sharing surface patterns (Figure 1).
@@
We explore the ability of large language models to detect and reason about cross-document coreference and hierarchy relationships, and present a novel method for this challenging setting. Given a concept mention in a scientific paper, our approach first automatically generates a concept definition: Definitions are generated by retrieving mention-specific literature and synthesizing a concept description that takes into account both global literature information and local context surrounding the mention. We augment original inputs with the definitions, and train LLMs to make use of them for detecting cross-document coreference and hierarchy. We further introduce relational definitions, which explicitly describe how two concept mentions are related or different. To avoid combinatorial explosion when generating relational definitions across all possible pairwise candidates, we create an efficient two-stage re-ranking approach that first generates singleton definitions and uses them to score and select candidates for which relational definitions will be computed.
@@
We train and evaluate our approach on the large and challenging SciCo dataset Cattan et al. (2021b), a benchmark for scientific cross-document coreference and hierarchy detection. In both fine-tuning (FT) and in-context learning (ICL) settings, our definition-based approach achieves significant improvements in detection accuracy. In the FT setting we obtain new state of art results, with singleton and relational definition augmentation leading to large gains in CoNLL F1 over models that don’t use definitions, and particularly strong gains in hierarchy detection. In the ICL setting, we evaluate the ability of GPT-4o-mini equipped with the advanced prompt optimizer DSPy Khattab et al. (2023); we find that when our definition augmentation is added, results substantially increase.
@@
Finally, we conclude with an extensive analysis of the strengths and limitations of contextualized, relational, and retrieval-based definitions, shedding light on when and how they contribute to performance. This analysis opens up intriguing avenues for future research, particularly in refining the generation and utilization of dynamic concept definitions and in improving the relational reasoning Alexander (2016) ability of LLMs over nuanced, fine-grained technical concepts.
@@ 2 Problem Setting
We focus on the fundamental problem of hierarchical cross document coreference resolution (H-CDCR), first introduced in Cattan et al. (2021b). In H-CDCR, our goal is to induce clusters of context-sensitive mentions that co-refer to the same concept, and to deduce a hierarchy among these concept clusters. Cattan et al. (2021b) created SciCo, a challenging dataset for training and evaluating H-CDCR models focusing on the complex scientific domain; in SciCo, the goal is to infer coreference clusters from sets of mentions extracted from computer science papers, and a hierarchy over the corefence clusters. For example, as illustrated in Figure 1, the goal would involve detecting that class imbalance and skewed label distribution belong to the same underlying concept cluster, that embedded networks does not belong in the same cluster as network embedding, and that network embedding belongs to a concept cluster that is a child of the parent concept representation learning.
Formally, we are provided with a collection of documents , each annotated with concept . Let be the set of mentions in document and the set of mentions across all . Similar to cross-document co-reference resolution, the first goal is to cluster the mentions in into disjoint clusters . Each cluster contains mentions that refer to the same underlying concept. The second goal is to determine a hierarchy over those clusters, specifically a hierarchical relation between cluster pairs that reflects on a ; a relation exists when the concept underlying entails a to . Given documents and mentions , the goal is to construct clusters and a hierarchy graph by learning from a set of N examples .
As we next discuss, we explore learning in both the fine-tuning (FT) and in-context learning (ICL) settings. Our novel approach involves augmenting mentions with dynamic, contextual definitions. We aim to synthesize definitions for specific mentions—and relational definitions explicitly describing one or more relations between specific pairs of mentions—such that the generated definitions assist models in more accurately inferring the correct cross-document relations.
@@ 3 Methods
Following Cattan et al. (2021b), we consider a multiclass classification setting where each mention pair can be assigned into four classes: (1) and co-refer (2) (3) and (0) none. Denote our definition augmentation model as a function , where is applied to a pair of candidate mentions and enriches each mention with a contextual definition. may correspond either to singelton definitions, in which case the definitions for and are generated in isolation of each other, or relational definitions in which the definition of depends on and vice versa. Equipped with that enriches mention contexts, during training we use a set of mentions to learn a pairwise scorer that minimizes
where represents one of the four possible classes, denotes the total number of training pairs, is a neural language model (§4 for details on specific LMs and input formatting).
@@
We embed all passages in using . Then, for every mention we retrieve a set of passages with maximal dot product similarity to . To reduce noise, we further follow common practice and apply a reranking and filtering step using a re-ranker LM that returns a new re-ranked and filtered context set , where is an ordered set of contexts for , after re-ranking and filtering (§4 for specific implementation details for ). Lastly, we feed into an LLM that generates contextual definition . In Figure 3 in the Appendix, we show the prompt used for generating ; see further implementation details in Section 4.
@@
More formally, for each mention and each candidate we use singleton definition augmentation to compute likelihood scores (each prediction of the model is a logit of size 4). We first filter out pairs for which the model predicted none (no relationship). Then, for each we compute the softmax probabilities and rank the predictions based on the model’s confidence in its assigned relation. We set a threshold value, , and for each mention , we generate relational definitions solely for the top pairs of mentions for which .
@@
Finally, given definitions we integrate them into LLMs for training. The current state of the art reported by Cattan et al. (2021b) was based on fine-tuning a Longformer model Beltagy et al. (2020). We fine-tune recent state of the art LLM, and also experiment with both few-shot prompting and the more advanced DSPy Khattab et al. (2023) prompt optimization framework (§ Appendix A.1).
@@ 4.1 Implementation Details
Generating definitions
As described earlier (§3) we used a RAG process for definition generation. For each definition generation iteration we retrieve the relevant papers, using mxbai-embed-large-v1 Lee et al. (2024); Li and Li (2023) as an embedding model with state-of-the-art performance. We use the Chroma DB333 https://github.com/chroma-core/chroma open-source vector database for fast retrieval using approximate nearest neighbor search. We rerank retrieved results with a cross-encoder (mxbai-rerank-large-v1444 mixedbread.ai/docs/reranking/models). Results are then finally fed into the definition generation model, for which we use Mixtral 8x7B, a mixture of experts LLM Jiang et al. (2024). We employ this method to produce singleton definitions for each individual mention and relational definitions (see Figures 3-4 for specific prompts).
@@
We train models both with the inclusion of definitions and without them. As base models, we report results for fine-tuning Mistral-7B555To check generalizablity of our results beyond Mistral, we also conduct a brief ablation experiment with another recent open source LLM Abdin et al. (2024). We observe similar overall trends. See Tables (10, 11, 12, 13, 14, 15) in appendix. Jiang et al. (2023), and Longformer Beltagy et al. (2020) which was used originally in Cattan et al. (2021b). We augment the original data to inject a definition for every mention. For Mistral-7B we construct a prompt (Figure 5) following the ChatML666 github.com/openai/openai-python/blob/release-v0.28.0/chatml.md format. For Longformer the input is processed as follows. For each mention and its corresponding paragraph, mention markers <m> and </m> surrounding the mention are used for obtaining a mention representation. We insert additional <def> and </def> markers surrounding each definition . Finally we concatenate all the above with the separation tokens <s> and </s> separating each mention (See Figure 6).
We fine-tuned the Mistral-7B model from the HuggingFace Wolf et al. (2019) library by adding a classification head, replacing the original head with a task-specific classification layer 777 huggingface.co/transformers/v3.0.2/model_doc/auto.html#automodelforsequenceclassification. The classification head is a simple feed-forward layer with a softmax activation function. Fine-tuning was conducted on 3 NVIDIA RTX A6000 48GB GPUs over 2 epochs. The training was carried out with a batch size of 2, 4 gradient accumulation steps and a maximum sequence length of 1664 tokens (these choices were made due to our academic resources). We used a default learning rate of 2e-5 with a weight decay of 0.001. We use Low-Rank Adaptation of LLM (LoRA), a parameter efficient fine-tuning method Hu et al. (2021) with the following parameters selected according to our resource constraints: lora_alpha: 16, lora_dropout: 0.1, lora_r: 8, SEQ_CLS as the task type and modules_to_save=["score"] in order to update and save the weights of the classification head we added to the model. We also attempted an alternative generative fine-tuning approach with the same setup, where the model was tasked with generating the correct relationship for each pair. However, this approach empirically led to worse results in our experiments.
@@
In our experiments with few-shot in-context learning (ICL) we use two types of prompts: one incorporating definitions and the other without them. We include one example per class (four total; DSPy further optimizes in-context examples). The prompt with definitions provides additional context for each example (see prompt in Figure 8), while the prompt without definitions relies solely on the examples (Figure 7). To account for variability in results due to few-shot selection, we report average results for 5 runs for each configuration, each time sampling a different set of few-shot examples and corresponding definitions (within each run, the same ICL examples are used for both with/without definition settings). We structure the output to extract classifications and reasoning automatically.888dspy-docs.vercel.app/docs/deep-dive/signature/executing-signatures#inspecting-output.
Due to the combinatorial nature of our task involving prediction of cross-document links across a potentially huge candidate space, we focus our ICL evaluation on a strong yet comparatively cheap model, GPT-4o-mini, and applied only to a subset of the test set (§4.2). Our ICL experiments cost roughly $500, while a stronger GPT-4o model would have incurred a cost of over $16000 which is beyond our academic budget. Concurrent results Lior et al. (2024) on SciCo with various open-source Mistral models in the few-shot setting led to very poor results in comparison to our results with GPT-4o-mini even without our definition-augmentation approach.
@@
In addition to standard co-reference resolution metrics (e.g., MUC, B³, CEAFe, and LEA), we employ the evaluation metrics proposed by Cattan et al. (2021b) for hierarchy evaluation. Specifically, we use the Cluster-Level Hierarchy Score, which evaluates the accuracy of hierarchical links between clusters, and the Path-Distance Score, which measures the degree of separation between concepts in the cluster graph. We refer the reader to Cattan et al. (2021b) for a detailed explanation of these metrics.
@@ 4.2 Results
| 1: … WAN expansion regardless of network structures and to enhance the network cyber-security… | 1: Protecting networks and information from unauthorized access … firewalls, antivirus software, encryption… | 1: Protecting networks … firewalls, encryption … In contrast to cyber-safety awareness, which is centered around individual behavior and education … |
@@
2: A range of operations involved in analyzing and understanding spoken language, including Speech Separation, Diarization, Automatic Speech Recognition … often serving as a crucial precursor to applications such as conversational speech synthesis …

### v3

@@ Abstract
We address the fundamental task of inferring cross-document coreference and hierarchy in scientific texts, which has important applications in knowledge graph construction, search, recommendation and discovery. Large Language Models (LLMs) can struggle when faced with many long-tail technical concepts with nuanced variations. We present a novel method which generates context-dependent definitions of concept mentions by retrieving full-text literature, and uses the definitions to enhance detection of cross-document relations. We further generate relational definitions, which describe how two concept mentions are related or different, and design an efficient re-ranking approach to address the combinatorial explosion involved in inferring links across papers. In both fine-tuning and in-context learning settings, we achieve large gains in performance on data subsets with high amount of different surfaces forms and ambiguity, that are challenging for models. We provide analysis of generated definitions, shedding light on the relational reasoning ability of LLMs over fine-grained scientific concepts.
@@ 1 Introduction
As the volume of scientific papers continues to explode, so does the intricate web of concepts mentioned across papers. Papers in computer science, for example, discuss fine-grained concepts such as algorithms, methods, tasks and problem areas Luan et al. (2018). Being able to automatically identify, e.g., when papers are referring to the same problem while using very different language (cross-document coreference), or that a method used in one paper is part of a more general family of methods (hierarchy)—could help unlock a wide range of applications, including in scientific knowledge-base construction Mondal et al. (2021); Hope et al. (2021), information retrieval Wang et al. (2023); Kang et al. (2024); Wu et al. (2024), recommendation systems Mysore et al. (2023), systems for paper reading and learning new topics Murthy et al. (2022); Guo et al. (2024), and even in computational creativity for science Hope et al. (2017); Portenoy et al. (2022); Wang et al. (2024); Srinivasan and Chan (2024).
@@
However, detecting cross-document coreference and hierarchy relations in scientific papers poses significant challenges Cattan et al. (2021b). Concepts manifest in long-tail, nuanced variations. Technical language is characterized by ambiguity, different surface forms and fine-grained levels of concept abstractions. As an example illustrating several of these challenges, consider energy-based network embedding, a specific method under the broader concept of neural representation learning despite no surface overlap or lexical cues that may suggest one concept subsumes the other; while low energy embedded networks in communications is unrelated, despite strong surface similarity (Figure 1). Surface forms network, energy, embedding/embedded refer to different scientific concepts in different contexts (ambiguity). Another example is class imbalance which refers to a skewed label distribution (Figure 1)—the same underlying scientific concept with different terms.
We explore the ability of large language models (LLMs) to detect cross-document coreference and hierarchy, i.e., detecting when scientific concept mentions across two papers co-refer (e.g., network embedding and neural representation of graphs), and when one concept mention is a parent/child of another mention (network embedding and the parent concept neural representation learning).
@@
We present a novel method for this challenging setting. Given a concept mention in a paper, our approach first generates a concept definition (e.g., network embedding is a neural representation of graphs…). Definitions are generated by retrieving literature relevant to the mention and synthesizing a concept description that takes into account both global literature information and the local context surrounding the mention. We augment original inputs with the definitions, and train LLMs to make use of them for detecting cross-document coreference and hierarchy. We further introduce relational definitions, which explicitly describe how two concept mentions are related or different. To avoid combinatorial explosion when generating relational definitions across all possible pairwise candidates, we create an efficient two-stage re-ranking approach that first generates singleton definitions and uses them to score and select candidates for which relational definitions will be computed.
@@
We train and evaluate our approach on the large and challenging SciCo dataset Cattan et al. (2021b), a benchmark for scientific cross-document coreference and hierarchy detection. The examples used above are real cases appearing in SciCo. In both fine-tuning (FT) and in-context learning (ICL) settings, our definition-based approach achieves substantial improvements in detection accuracy, particularly on challenging data subsets marked by a high amount of different surface forms and ambiguity. These hard cases account for 10–20 percent of the test data. In the FT setting, we obtain new state-of-the-art results, with singleton and relational definition augmentation leading to large gains in CoNLL F1 over models that do not use definitions, and particularly strong gains in hierarchy detection. In the ICL setting, we evaluate the ability of GPT-4o-mini equipped with the advanced prompt optimizer DSPy Khattab et al. (2024); we find that when our definition augmentation is added, results substantially increase.
@@
Finally, we conclude with an extensive analysis of the strengths and limitations of contextualized, relational, and retrieval-based definitions, shedding light on when and how they contribute to performance. This analysis opens up intriguing avenues for future research, particularly in refining the generation and utilization of dynamic concept definitions and in improving the relational reasoning Alexander (2016) ability of LLMs over nuanced, fine-grained technical concepts.111Code, models and generated definitions available at https://github.com/TomHopeLab/ScicoRadar.
@@ 2 Problem Setting
We focus on the fundamental problem of hierarchical cross document coreference resolution (H-CDCR), first introduced in Cattan et al. (2021b). In H-CDCR, our goal is to induce clusters of context-sensitive mentions that co-refer to the same concept, and to deduce a hierarchy among these concept clusters. Cattan et al. (2021b) created SciCo, a challenging dataset for training and evaluating H-CDCR models focusing on the complex scientific domain; in SciCo, the goal is to infer coreference clusters from sets of mentions extracted from computer science papers, and a hierarchy over the co-reference clusters. For example, as illustrated in Figure 1, the goal would involve detecting that class imbalance and skewed label distribution belong to the same underlying concept cluster, that embedded networks does not belong in the same cluster as network embedding, and that network embedding belongs to a concept cluster that is a child of the parent concept representation learning.
Formally, we are provided with a collection of documents , each annotated with concept . Let be the set of mentions in document and the set of mentions across all . Similar to cross-document co-reference resolution, the first goal is to cluster the mentions in into disjoint clusters . Each cluster should contain mentions that refer to the same underlying concept. The second goal is to determine a hierarchy over those clusters, specifically a hierarchical relation between cluster pairs that reflects on a ; a relation exists when the concept underlying entails a to . Given documents and mentions , the goal is to construct clusters and a hierarchy graph by learning from a set of N examples .
As we discuss next, we explore learning in both the fine-tuning (FT) and in-context learning (ICL) settings. Our novel approach involves augmenting mentions with dynamic, contextual definitions. We aim to synthesize definitions for specific mentions—and relational definitions explicitly describing one or more relations between specific pairs of mentions—such that the generated definitions assist models in more accurately inferring the correct cross-document relations.
@@ 3 Methods
Following Cattan et al. (2021b), we consider a multiclass classification setting where each mention pair can be assigned into four classes: (1) and co-refer (2) (3) and (0) none. We denote our definition augmentation model as a function , where is applied to a pair of candidate mentions and enriches each mention with a contextual definition. may correspond either to singelton definitions, in which case the definitions for and are generated in isolation of each other, or relational definitions in which the definition of depends on and vice versa. Equipped with that enriches mention contexts, during training we use a set of mentions to learn a pairwise scorer Yang et al. (2022) that minimizes
where represents one of the four possible classes, denotes the total number of training pairs, is a neural language model.
@@
We embed all passages in using . Then, for every mention we retrieve a set of passages with maximal dot product similarity to . To reduce noise, we further follow common practice and apply a reranking and filtering step using a re-ranker LM that returns a new re-ranked and filtered context set , where is an ordered set of contexts for , after re-ranking and filtering. Lastly, we feed into an LLM that generates contextual definition . In Figure 3, we show the prompt logic used for generating (see technical details in Appendix A.1).
@@
More formally, for each mention and each candidate we use singleton definition augmentation to compute likelihood scores (each prediction of the model is a logit of size 4). We first filter out pairs for which the model predicted none (no relationship). Then, for each we compute the softmax probabilities and rank the predictions based on the model’s confidence in its assigned relation. We set a threshold value to select the top 25% pairs as a tradeoff between accuracy and computational efficiency; we compare different values of this threshold on the Hard-10 subset in Appendix A.4). For each mention , we generate relational definitions solely for the top pairs of mentions for which .
@@
Finally, given definitions we integrate them into LLMs for training. The current state of the art reported by Cattan et al. (2021b) was based on fine-tuning a Longformer model Beltagy et al. (2020). We fine-tune recent state of the art LLM, and also experiment with both few-shot prompting and the more advanced DSPy Khattab et al. (2024) prompt optimization framework (see more details in Appendix A.3).
@@ 4 Experiments
In this section we describe our experimental setup and results. We first outline the implementation details (§4.1), and then present the main results (§4.2), which introduce the baselines and report performance on both the hardest subsets and the overall dataset.
@@ 4.1 Implementation Details
We train models both with the inclusion of definitions and without them. As base models, we report results for fine-tuning Mistral-7B333To check generalizablity of our results beyond Mistral, we also conduct a brief ablation experiment with another recent open source LLM Abdin et al. (2024). We observe similar overall trends. See Tables 14–19 in Appendix. Jiang et al. (2023), and Longformer Beltagy et al. (2020) which was used originally in Cattan et al. (2021b). We augment the original data to inject a definition for every mention. For Mistral-7B we construct a prompt (Figure 6) following the ChatML444github.com/openai/openai-python/blob/release-v0.28.0/chatml.md format. For Longformer the input is processed as follows. For each mention and its corresponding paragraph, mention markers <m> and </m> surrounding the mention are used for obtaining a mention representation. We insert additional <def> and </def> markers surrounding each definition . Finally we concatenate all the above with the separation tokens <s> and </s> separating each mention (See prompt in Figure 7).
We fine-tuned the Mistral-7B model from the HuggingFace Wolf et al. (2020) library by adding a classification head, replacing the original head with a task-specific classification layer.555huggingface.co/transformers/v3.0.2/model_doc/auto.html#automodelforsequenceclassification The classification head is a simple feed-forward layer with a softmax activation function (see Appendix A.2 for hyper parameters, hardware and implementation details). We also attempted an alternative generative fine-tuning approach with the same setup, where the model was tasked with generating the correct relationship for each pair. However, this approach led to worse results in our experiments.
@@
In our experiments with few-shot in-context learning (ICL) we use two types of prompts: one incorporating definitions and the other without them. We include one example per class (four total; DSPy further optimizes in-context examples). The prompt with definitions provides additional context for each example (see prompt in Figure 9), while the prompt without definitions relies solely on the examples (Figure 8). To account for variability in results due to few-shot selection, we report average results for 5 runs for each configuration, each time sampling a different set of few-shot examples and corresponding definitions (within each run, the same ICL examples are used for both with/without definition settings). We structure the output to extract classifications and reasoning automatically.666dspy-docs.vercel.app/docs/deep-dive/signature/executing-signatures#inspecting-output.
Due to the combinatorial nature of our task with cross-document links across a potentially huge candidate space, we focus ICL evaluation on a strong yet comparatively cheap model, GPT-4o-mini, and applied only to a subset of the test set (§4.2). Our ICL experiments cost roughly $500, while a stronger GPT-4o model would have incurred a cost of over $16000 which is beyond our academic budget. Concurrent results Lior et al. (2024) on SciCo with various open-source Mistral models in the few-shot setting led to very poor results in comparison to our results with GPT-4o-mini even without our definition-augmentation approach.
@@
For co-reference evaluation, we adopt the standard metrics MUC, B3, CEAFe, and LEA, and report the widely used CoNLL F1, defined as the average of the MUC, B³, and CEAFe F1 scores, as our primary measure. To evaluate the hierarchy, we use the following evaluation methods created in Cattan et al. (2021b):
Cluster-Level Hierarchy Score checks whether each predicted parent–child link between mention clusters aligns with at least one mention pair in the gold hierarchy. This design avoids “double penalizing” minor co-reference mismatches, For example, if the system predicts a parent–child relationship between clusters [Information Extraction] → [Definition Extraction, Pattern-based Extraction], and at least one pair of mentions (e.g., “Information Extraction” → “Definition Extraction”) exists in the gold hierarchy, this counts as correct. We also consider a stricter version (50% Hierarchy Score) that requires at least half of the mentions in the predicted cluster pair to align with the gold parent–child link.
Path-Distance Score compares distances between mentions (in terms of the number of edges separating them) in the predicted hierarchy vs. the gold. It provides partial credit when the distances are close, even if not perfectly matched. For instance, consider the mentions “Information Extraction” and “Pattern-based Definition Extraction”: if they are separated by two edges in the gold hierarchy but only one edge in the predicted hierarchy, the metric gives partial credit proportional to the closeness of these path lengths.
We refer the reader to Cattan et al. (2021b) for full technical details on these hierarchy metrics.
@@ 4.2 Results
We next present the experimental results. We begin by establishing baselines and evaluating on the hardest subsets of the dataset, where the task is most challenging and the benefits of our definition based approach has the most impact. We then assess performance on the full dataset to measure overall improvements, and conclude with an ablation study that examines the specific contribution of different types of definitions.
| Metric | SciCo-Longformer (Baseline) | SciCo-LongformerSingleton def | ||||
| All | Hard-20 | Hard-10 | All | Hard-20 | Hard-10 | |
| CoNLL F1 | 76.46 | 68.18 | 64.34 | 78.18 | 71.10 | 65.28 |
| Hierarchy F1 | 46.09 | 46.44 | 45.46 | 47.43 | 49.28 | 46.21 |
| Hierarchy F1 50% | 35.14 | 29.61 | 28.07 | 38.52 | 33.77 | 29.52 |
| Path Ratio | 44.91 | 36.21 | 34.26 | 48.81 | 44.07 | 35.36 |
Evaluation focus
In addition to reporting overall performance (see Table 3), we focus on the “hard” data subsets provided by Cattan et al. (2021b), comprised of the 10% (Hard-10) or 20% (Hard-20) topics (collections of candidate mentions) with lowest scores as ranked according to CoNLL F1 by using Levenshtein distance in agglomerative clustering as a baseline. The subsets are particularly challenging, with a high amount of different surface forms and ambiguity, and are thus of most interest; e.g., the best model reported by Cattan et al. (2021b), achieved only 64.34 F1 for coreference CoNLL F1 on the Hard-10 subset, while achieving 76.46 F1 on the full test set.
| Model | Coreference | Hierarchy | Path | |
| CoNLL F1 | F1 | F1-50% | Ratio | |
| SciCo-Longformer (Baseline) | 76.46 | 46.09 | 35.14 | 44.91 |
| SciCo-Mistral-7B | 81.82 | 56.84 | 49.70 | 57.21 |
| SciCo-RadarSingleton | 81.95 | 57.33 | 50.58 | 57.27 |
| SciCo-RadarRelational def | 82.07 | 58.03 | 50.72 | 57.60 |
Baselines
As a primary baseline, we use the model introduced in the original SciCo paper Cattan et al. (2021b), which defines the standard evaluation setting for this task. In addition, we include our own fine-tuned models that do not use definitions, in order to isolate the contribution of definition based augmentation. We also report results for in-context learning (ICL) with GPT-4o-mini, to establish a baseline for definition augmentation without finetuning. While SEAM Lior et al. (2024) also reports results on SciCo and uses more recent models, its performance falls well below that of the original SciCo baseline.
Evaluation summary
We report our results in Table 1 (Hard-10), Table 2 (Longformer all data) and Table 3 (all data) from the main paper, along with the full Tables [14, 15, 16, 17, 18, 19] in the Appendix (results in all subsets reflect similar model trends). ICL few-shot results are averaged across 5 runs as described earlier. We first observe that, not very surprisingly, fine-tuning of recent LLMs leads to large gains over the smaller Longformer model used originally in Cattan et al. (2021b). Fine-tuning also achieve considerable gains over concurrent results for SciCo CDCR in the few-shot setting with larger variants from the Mistral family (see Table 3 in Lior et al. (2024)), highlighting the value of fine-tuning in this task. More interestingly, our contextualized definitions led to strong improvements, for both LLM fine tuning and ICL. Our SciCo-RadarRelational model, which consists of fine-tuning Mistral-7B on SciCo augmented by relational definitions, achieved the highest results in every evaluation category, as exemplified in Table 1. Generally, relational definitions outperformed singleton definitions, by carrying more nuanced information about relationships that the model might miss with the more generic singleton definitions777Paired bootstrap tests (Appendix A.5) confirm statistically reliable gains for hierarchy metrics across both hard subsets and for CoNLL F1 on Hard-10, with the remaining metrics reflecting positive upward trends. (see Section 5 for more analysis). ICL with larger LLMs OpenAI (2024), including prompt optimization with DSPy, underperformed smaller fine-tuned models in our experiments, though future work may explore this further with more models and in-context examples.
As seen in Table 2, augmenting Longformer with our contextual definitions also resulted in substantial enhancements in both hierarchy and coreference scores, by up to 3.38 and 1.72 respectively. Even more substantial improvements were observed within the Hard-20 subsets, particularly in hierarchy and hierarchy 50% scores, achieving improvements of 2.84 and 4.16. With Mistral-7B, incorporating definitions consistently improved all metrics. As seen In Table 1, incorporating definitions notably boosted Mistral’s hierarchy F1 and F1-50% scores by up to 4.5 and 6.1 using singleton definitions, and by 6.78 and 8.63 with relational definitions.
We also report additional coreference and hierarchy precision-recall metrics in Tables 14, 15 and 17, 18 in the Appendix. For coreference, the F1 gains with relational definitions come from a substantial boost in recall with only a relatively small dip in precision. For hierarchy, we see a similar trend where both singleton and relational definitions improve recall while maintaining stable precision.
Finally, to assess the stability of ICL results with random few-shot examples or examples optimized by random search in DSPy, we report average results for five runs (Table 4).888The SciCo dataset contains 69,682 mention pairs in the Hard-10 subset, along with extensive context and definitions per mention; running our ICL experiments on this subset cost $500. The full test set contains 564,352 mention pairs. We find that the effects are overall stable, with definitions helping boost results across both the few-shot and DSPy experiments;999We also experimented with several runs of fine-tuning, finding negligible variance across seeds. interestingly, we do not observe gains from using DSPy in our experiments.
| Model | Coreference | Hierarchy | Path | |
| CoNLL F1 | F1 | F1-50% | Ratio | |
| GPT 4o-mini | 51.28 ± 2.21 | 40.22 ± 1.70 | 16.95 ± 1.62 | 31.04 ± 2.23 |
| GPT 4o-miniDSPy | 50.71 ± 0.66 | 42.23 ± 0.94 | 14.74 ± 1.36 | 27.41 ± 1.29 |
Overall vs. Hard Subsets
Table 3 shows only modest overall improvements on the full SciCo test set, whereas Table 1 reveals much larger gains on the Hard-10 and Hard-20 subsets. This is expected, since many mentions can be resolved by simple lexical overlap or the model’s prior knowledge, reducing the benefits of adding definitions. By contrast, the Hard-10/Hard-20 cases feature ambiguous concepts and more variations of different surface forms, precisely where retrieval-based definitions can clarify those subtle distinctions or shared properties, thus offering substantial improvements. Additionally, these challenging cases make up only 10–20% of the data, so the strong gains do not dramatically affect the overall average. Nonetheless, linking these rare or nuanced mentions accurately is crucial for downstream scientific tasks, for example, knowledge graph construction or document retrieval, where mislinking can lead to incorrect representations of fine-grained concepts. A similar phenomenon was observed by Ravi et al. (2023) in event coreference, where the majority of mention pairs were straightforwardly matched by a model’s surface or parametric knowledge. Substantial benefits from injecting external knowledge was reported primarily for the smaller more challenging fraction of mentions, in line with our findings on the Hard-10/Hard-20 subsets. Notably, in table 2, we observe that the gains from adding definitions for the Longformer model are larger on the full and the Hard-20 test sets compared to the Hard-10 subset, which contrasts with the trends we observed for stronger models. We hypothesize that Longformer struggles more with the inherently ambiguous and challenging examples in the Hard 10 subset, limiting its ability to fully exploit the additional context provided by definitions. Consequently, while definition augmentation still improves performance, the relative gains are smaller on the hardest cases compared to simpler ones.
@@
| 1: … WAN expansion regardless of network structures and to enhance the network cyber-security… | 1: Protecting networks and information from unauthorized access … firewalls, antivirus software, encryption… | 1: Protecting networks … firewalls, encryption … In contrast to cyber-safety awareness individual behavior and education … |
@@
2: Operations involved in analyzing and understanding spoken language, including Speech Separation, Diarization, Automatic Speech Recognition … a precursor to applications such as conversational speech synthesis …